OpenStudy (yb1996):

A particle moves on the curve y = 3e^(x^2 + 1). Find the rate of change of the y-coordinate at the instant the x-coordinate is 2 ad the x-coordinate is changing at 1/12 units/second. Can someone please post the solution to this problem. I will give a medal.

OpenStudy (freckles):

looks like you are suppose to take x' as 1/12 instead of 1

OpenStudy (freckles):

So before we consider x' as something other than 1 do you know how to find dy/dx?

OpenStudy (yb1996):

dy/dx would just be 6xe^(x^2+1), right?

OpenStudy (freckles):

\[\text{ so } y'=3e^{x^2+1}(2x)x' \\y'=6xe^{x^2+1}x' \\ y'=6xe^{x^2+1} \cdot \frac{1}{12} \\ y'=\frac{1}{2} xe^{x^2+1}\] Find y' at x=2

OpenStudy (yb1996):

On the first line you wrote that \[y' = 3e ^{x^2+1} (2x) x'\], how did you get x' in the equation?

OpenStudy (freckles):

x' is the rate of x

OpenStudy (freckles):

x' isn't 1 here

OpenStudy (freckles):

x' is 1/12

OpenStudy (yb1996):

ok, apparently the derivative that I originally got did not have x' in it, can you please explain to me what I had done wrong?

OpenStudy (freckles):

\[D(x,y)=\sqrt{x^2+y^2} \\ D^2=x^2+y^2 \\ \text{ by chain rule we have } \\ 2D \frac{dD}{dt}=2x \frac{dx}{dt}+2y \frac{dy}{dt} \\ D \frac{dD}{dt}=x \frac{dx}{dt}+2 y \frac{dy}{dt}\] and we just found dy/dt

OpenStudy (freckles):

dx/dt is given

OpenStudy (freckles):

oops that equation shouldn't have that 2 in it

OpenStudy (yb1996):

so we are taking the derivative in respect to t?

OpenStudy (freckles):

\[D(x,y)=\sqrt{x^2+y^2} \\ D^2=x^2+y^2 \\ \text{ by chain rule we have } \\ 2D\frac{dD}{dt}=2x \frac{dx}{dt}+2y \frac{dy}{dt} \\ D \frac{dD}{dt}=x \frac{dx}{dt}+y \frac{dy}{dt}\]

OpenStudy (freckles):

yep

OpenStudy (freckles):

x'=dx/dt and y'=dy/dt

OpenStudy (freckles):

there has to be another variable in which x depends since x has rate other than 1

OpenStudy (yb1996):

ahhhhh, that means that we have to do implicit differentiation in this problem, that makes sense. Thanks for the help!